OpenStudy (anonymous):

Solve the equation. sin theta = -0.3 It's supposed to have more than one answer. Its solving basic trig equations.

OpenStudy (anonymous):

162.5

OpenStudy (johnweldon1993):

All you must to to solve for theta is take the arcsin of both sides of the function \[\large \sin^{-1}\sin(\theta) = \sin^{-1}(-0.3)\] \[\large \theta = \sin^{-1}0.3\] \[\large \theta = -17.46^\circ\]